Suppose that a drop of mist is a perfect sphere and that, through condensation, the drop picks

up moisture at a rate proportional to its surface area. Under these circumstances, what will be

the e‚ect on the drop’s radius.



1
Expert's answer
2021-10-29T03:01:13-0400

Solution:

"V=\dfrac43\pi r^3
\\ \dfrac{dV}{dt}=\dfrac43.3\pi r^2 \dfrac{dr}{dt}
\\=4\pi r^2 \dfrac{dr}{dt}\ ...(i)"

"A=4\pi r^2 ...(ii)"

It is given that "\dfrac{dV}{dt}=k.(Surface\ area)"

"\Rightarrow 4\pi r^2 \dfrac{dr}{dt}=k.(4\pi r^2)" [using (i) and (ii)]

"\Rightarrow \dfrac{dr}{dt}=k"

It shows radius increases at a constant rate.